A particle starts from rest and travels in a straight line so that t seconds after passing a fixed point O,
its acceleration a m/S^2 is given by a=2(3-e^-t)
Calculate
(a)the velocity of the particle when t=2
(b)the displacement of the particle from O when t=1

Found 2 solutions by rothauserc, ikleyn:
Answer by rothauserc(4718)

Refer to kinematics of motion equations
:
(a) Since the acceleration is constant and Vi = 0
:
Vf = at
:
Vf = 2(3-e^-2) * 2 = 4(3-e^-2) m/sec
:
(b) Vf = 2(3-e^-2) * 1 = 2(3-e^-2)
:
s = (Vi + Vf)/2 * t
:
s = 2(3-e^-2)/2 * 1 = (3-e^-2) m
:

Answer by ikleyn(52864)

.
A particle starts from rest and travels in a straight line so that t seconds after passing a fixed point O,
its acceleration a m/S^2 is given by a=2(3-e^-t)
Calculate
(a)the velocity of the particle when t=2
(b)the displacement of the particle from O when t=1
~~~~~~~~~~~~~~~~


The post by @rothauserc consists of two parts, and both parts are wrong.


The first part is wrong, since @rothausers MISTAKENLY assumes that the acceleration "a" is constant,
on contrary to the given input info.

The second part is wrong because the average speed is half of the sum
of the initial and final speed ONLY for UNIFORMLY accelerated move,

which IS NOT the case.